# Potential Pitfalls in Visual Models of Tipping Points - And How to Fix Them

### v1

@@ 5 Tipping Dynamics and Popular Visualizations: From Cell Dynamics to Climate System Tipping Points
In the following, we argue that both the ensemble interpretation and the single system (i.e. probability) interpretation are important in real systems and must be carefully distinguished depending on the application. We discuss examples from literature regarding this distinction between deterministic/nondeterministic processes on the one hand, and the ensemble/probability interpretation on the other hand, and show that these subtle distinctions are not carefully made in all cases.
@@ 5.1 Tipping Points in Cell Dynamics
5.2 Climate Systems Tipping Points
@@
Climate systems, such as the Amazon rainforest, the Greenland and Antarctic ice sheets, various monsoons or boreal permafrost, are complex dynamical systems that usually reside in a stable equilibrium determined by internal processes and boundary conditions. Changing external conditions, such as an increase in greenhouse gas emissions or rising global surface temperatures, can lead to qualitative changes in the stability landscape of the system, resulting in B-tipping. Tipping of climate systems can also be caused by stochastic forcing of the system state (N-Tipping), such as weather patterns and extreme weather events, and by human intervention in ecosystems. These are expected to have an even greater destabilizing effect on systems as global temperatures rise and the stability and resilience of climate systems decreases (see Interaction of B- and N-tipping in section 3). Stochasticity can also be used to model the uncertainty introduced by the complexity of the physical processes involved in these climate systems (see section 3). However, this uncertainty about the current and future state of the system can be difficult to incorporate into visualizations. Although the purpose of visualizations of the type shown in Fig. 7, taken from Steffen et al. 2018 [1], is clear, several points can be misunderstood at the level of the individual mental model induced by the visualizations.
@@ 6 Applications: From Cell Differentiation to Tipping of the Amazon Rainforest
As example systems, we introduce two simple one-variable models of cell differentiation and tipping behavior in the Amazon rainforest, respectively. We present different visualization methods to incorporate uncertainty and stochasticity into the potential picture, according to the system type - ensemble or single system.
@@ 6.1 Cell Differentiation
To mathematically model the process of cell differentiation, we adapt a simplified model from Ferrell, 2012 [5] that models cell fate induction, the process by which cells commit to a particular lineage or function through signals from gene regulatory networks or neighboring cells. In our adapted model, the variable $x$ represents the synthesis of a differentiation regulator, the molecule responsible for promoting cell differentiation. These regulators are usually proteins such as transcription factors or signaling molecules that help ”guide” the cell toward a particular fate and stimulate its own production. Their rate of synthesis is given by a baseline rate $\eta$ (the regulator is constantly produced at a certain rate), which serves as a bifurcation parameter. As more of the protein is present in the cell, even more of it is produced - a positive feedback loop which is modeled by a Hill function [34]. The total synthesis rate is then given by $\eta+\alpha x^{5}/(K^{5}+x^{5})$, where the parameter $K$ represents the concentration of $x$ at which the response rate reaches half its maximum. $K$ essentially acts like a threshold - when $x$ approaches $K$, the feedback loop is strongly engaged. Additionally, the regulator molecule is constantly being broken down or removed from the cell at the so-called degradation rate $\beta x$, which is assumed to be linear. Combining these different terms results in the differential equation
@@
For simplicity, we follow [5] in assuming that the variables are dimensionless. With the parameter values $K=1$, $\alpha=1$, and $\beta=0.5$, the system exhibits a saddle-node bifurcation at $(x_{b},\eta_{b})\approx(0.58,0.23)$.
We assume that an inductive stimulus, such as an external signal like a hormone or interaction with other cells, leads to a slow increase in the baseline synthesis rate $\eta$, starting from a small initial value of $\eta=0.05$. When the synthesis rate of the regulator reaches and exceeds the threshold value $\eta\approx 0.23$, the undifferentiated state of the cell disappears at $x\approx 0.58$, forcing the cell to adopt the differentiated state at $x\approx 2.44$. This means that the cell suddenly has to commit to a certain fate - a tipping process.
Adding an additive noise term to the equation (10) now allows us to model slight variations in the synthesis $x$ of the differentiation regulator caused by external influences or chemical/biological variations in the cell’s environment. An ensemble of undifferentiated cells could then commit to a differentiated state at different stimulus intensities - or not at all.
@@
In the case of B-tipping (Fig. 9a), all cells of the ensemble eventually transition to the differentiated state after the system crosses the threshold (i.e. bifurcation point), while some systems start transitioning even before due to stochastic forcing. This can also be seen for the case without B-tipping (Fig. 9b), as even though the deterministic threshold is not crossed, noise forces a small number of systems over the potential threshold (forcing cells to commit to a fate).
@@ 6.2 Amazon Rainforest
[…]
@@
[…]
@@
[…]
@@
Special care must be taken when interpreting this visualization. Climate system dynamics are obviously driven by deterministic physical processes, and stochasticity is introduced only to model internal and external uncertainties and influences of the system. However, it is generally impossible to determine the exact state of the system and its distance to the tipping point, so the exact tipping point boundary can usually only be determined after the system has tipped - an outcome to be avoided in the context of climate dynamics. That is, if the model/visualization shows that the system state has transitioned to a tipped state with some non-zero probability, it is usually easy to validate whether the system has actually tipped. However, in the context of climate systems, the timescales for tipping to occur after the threshold is reached can vary from years to centuries [2]. Thus, an increase in the probability of the tipped state within the model signals that the real world system may be approaching a tipping point and that the probability of the system tipping in the future is increasing. Given that there is only one Earth, it is ”too risky to bet against’ [41].
@@ 7 Summary and Outlook
In this paper, we critically discussed potential pitfalls in visualizations of nonlinear dynamics used in both science communication and research. We discussed the need to distinguish between the purpose of a visualization, as intended by the author, and the individual mental model induced in the recipient, which also depends e.g. on his prior knowledge. We also argued that, in general, mere analogies (such as the ”potential landscape” in classical mechanics compared to that in nonlinear dynamics, see Fig. 2) must be carefully distinguished from mathematically sound graphical representations. In nonlinear dynamics we have to distinguish between deterministic and stochastic models, and between ensembles and single systems. Using examples representative of visualizations in popular and academic literature, we have shown that these distinctions are not always made. Moreover, we argued that the introduction of quasi-potentials for the illustration of stochastic processes might be misleading as discussed in Sect. 3. We therefore propose the visualizations shown in Fig. 5 for nonlinear dynamics, carefully distinguishing between ensemble interpretation and individual system interpretation. As explicit applications, we discussed a model of cell dynamics (Fig. 9, ensemble interpretation) and a model of the Amazon rainforest (Fig. 11, individual system). These visualization techniques can be generalized to any other nonlinear system, and we believe that they could help to clearly distinguish the different cases discussed in 1. In future work, empirical research will be important in order to test whether the individual mental models induced match with the intended purpose of these visualization.

### v2

@@ 5 Tipping Dynamics and Popular Visualizations: From Cell Dynamics to Climate System Tipping Points
In the following, we argue that both the ensemble interpretation and the single system (i.e., probability) interpretation are important in real systems and must be carefully distinguished depending on the application. We discuss examples from literature regarding this distinction between deterministic/nondeterministic processes on the one hand, and the ensemble/probability interpretation on the other hand, and show that these subtle distinctions are not carefully made in all cases.
@@ 5.1 Tipping points in cell dynamics
5.2 Tipping points in climate systems
@@
Climate systems, such as the Amazon rainforest, the Greenland and Antarctic ice sheets, various monsoons or boreal permafrost, are complex dynamical systems that usually reside in a stable equilibrium determined by internal processes and boundary conditions. Changing external conditions, such as an increase in greenhouse gas emissions or rising global surface temperatures, can lead to qualitative changes in the stability landscape of the system, resulting in B-tipping. Tipping of climate systems can also be caused by stochastic forcing of the system state (N-Tipping), such as weather patterns and extreme weather events, and by human intervention in ecosystems. These are expected to have an even greater destabilizing effect on systems as global temperatures rise and the stability and resilience of climate systems decreases (see Interaction of B- and N-tipping in section 3). Stochasticity can also be used to model the uncertainty introduced by the complexity of the physical processes involved in these climate systems (see section 3). However, this uncertainty about the current and future state of the system can be difficult to incorporate into visualizations. Although the purpose of visualizations of the type shown in Fig. 7, taken from Steffen et al. 2018 [37], is clear, several points could be misunderstood at the level of the individual mental model that might be induced by the visualizations.
@@ 6 Applications: From Cell Differentiation to Tipping of the Amazon Rainforest
To demonstrate how our proposed visualization framework can be applied to concrete physical scenarios, we now ground our discussion in two simple, established, one-variable models of system tipping. The first, from molecular biology, describes cell differentiation, an example of an ensemble of systems. The second, from climate science, models the potential dieback of the Amazon rainforest, an example of a single system. We present visualization methods incorporating uncertainty and stochasticity into the potential picture of the system, illustrating the practical importance of the conceptual distinctions made in this paper.
@@ 6.1 Cell differentiation
To mathematically model the process of cell differentiation, we adapt a simplified model from Ferrell and James [11] that models cell fate induction, the process by which cells commit to a particular lineage or function through signals from gene regulatory networks or neighboring cells. In our adapted model, the variable $x$ represents the synthesis of a differentiation regulator, the molecule responsible for promoting cell differentiation. These regulators are usually proteins such as transcription factors or signaling molecules that help ”guide” the cell toward a particular fate and stimulate its own production. Their rate of synthesis is given by a baseline rate $\eta$ (the regulator is constantly produced at a certain rate), which serves as a bifurcation parameter. As more of the protein is present in the cell, even more of it is produced - a positive feedback loop which is modeled by a Hill function [14]. The total synthesis rate is then given by $\eta+\alpha\frac{x^{5}}{K^{5}+x^{5}}$, where the parameter $K$ represents the concentration of $x$ at which the response rate reaches half its maximum. $K$ essentially acts like a threshold - when $x$ approaches $K$, the feedback loop is strongly engaged. Additionally, the regulator molecule is constantly being broken down or removed from the cell at the so-called degradation rate $\beta x$, which is assumed to be linear. Combining these different terms results in the differential equation
@@
For simplicity, we follow Ferrell and James [11] in assuming that the variables are dimensionless. With the parameter values $K=1$, $\alpha=1$, and $\beta=0.5$, the system exhibits a saddle-node bifurcation at $(x_{b},\eta_{b})\approx(0.58,0.23)$. We assume that an inductive stimulus, such as an external signal like a hormone or interaction with other cells, leads to a slow increase in the baseline synthesis rate $\eta$, starting from a small initial value of $\eta=0.05$. When the synthesis rate of the regulator reaches and exceeds the threshold value $\eta\approx 0.23$, the undifferentiated state of the cell disappears at $x\approx 0.58$, forcing the cell to adopt the differentiated state at $x\approx 2.44$. This means that the cell suddenly has to commit to a certain fate - a tipping process.
Combining an additive noise term to the equation (11) now allows us to model slight variations in the synthesis $x$ of the differentiation regulator caused by external influences or chemical/biological variations in the cell’s environment. An ensemble of undifferentiated cells could then commit to a differentiated state at different stimulus intensities - or not at all.
@@
In the case of B-tipping (Fig. 9a), all cells of the ensemble eventually transition to the differentiated state after the system crosses the threshold (i.e., bifurcation point), while some systems start transitioning even before due to stochastic forcing. This can also be seen for the case without B-tipping (Fig. 9b), as even though the deterministic threshold is not crossed, noise forces a small number of systems over the potential threshold (forcing cells to commit to a fate).
@@ 6.2 Amazon rainforest
Here, we model the Amazon rainforest using a simplified version of a TRIFFID model (top-down representation of interactive foliage and flora including dynamics) [8] for a single vegetation type, with modifications and parameter values taken from Ritchie et al. [29]. Its dependent variable $v$ represents the proportion of vegetation in the Amazon ecosystem and is modeled using a logistic equation with an additional decay term
$\frac{\text{d}v}{\text{d}t}=gv(1-v)-\gamma v.$ (12)
Here, $\gamma=0.2/\text{yr}$ is the disturbance rate representing vegetation mortality and $g$ is the growth rate, a function of the local temperature $T_{l}$, given by
$g=g_{0}\left[1-\left(\frac{T_{l}-T_{\text{opt}}}{\beta}\right)^{2}\right]$ (13)
The maximum growth rate $g_{0}=2/\text{yr}$ is reached when the local temperature $T_{l}$ is equal to the optimal growth temperature $T_{\text{opt}}=28\text{\textdegree C}$ and $\beta=10\text{\textdegree C}$ determines the half-width of the growth-temperature curve. Finally, there is an additional feedback between the vegetation fraction $v$ and the local temperature; an increase in vegetation cover leads to a lower local temperature $T_{l}=T_{f}+\alpha\,(1-v)$. The constant $\alpha=5\text{\textdegree C}$ represents the difference in surface temperature between bare soil $(v=0)$ and complete forest cover $(v=1)$, and $T_{f}$ represents the local temperature if the region were completely covered by vegetation, which depends only on the global temperature and is therefore used as an external forcing parameter. The final differential equation is then given by
@@
Starting at $T_{f}\approx 33.5\text{\textdegree C}$, $T_{f}$ increases with rising global temperatures, up to a saddle-node bifurcation at $(T_{f,b},v_{b})\approx(34.7\text{\textdegree C},0.66)$, where the stable equilibrium of a partially vegetated Amazon at $v\approx 0.66$ vanishes, causing the system to tip to the remaining stable equilibrium at $v=0$, an Amazon without vegetation cover. Adding noise to the system allows us to model real-world perturbative influences such as weather effects and human intervention.
@@
Figure 11: Potential visualization of two different single system trajectories of a tipping process representing dieback of the Amazon rain forest. The bifurcation parameter dynamics (Indirect measure of global temperature) are as detailed in Fig. 10, starting at $33.5\text{\textdegree C}$, eventually (a) crossing the tipping threshold ($T_{f}\approx 34.7\text{\textdegree C}$) and (b) approaching, but not crossing it. Time increases along the $z$-axis towards the viewer.
@@
Special care must be taken when interpreting this visualization. Climate system dynamics are obviously driven by deterministic physical processes, and stochasticity is introduced only to model internal and external uncertainties and influences of the system. However, it is generally impossible to determine the exact state of the system and its distance to the tipping point, so the exact tipping point boundary can usually only be determined after the system has tipped - an outcome to be avoided in the context of climate dynamics. That is, if the model/visualization shows that the system state has transitioned to a tipped state with some non-zero probability, it is in principle possible to validate whether the system has actually tipped. However, in the context of climate systems, the timescales for tipping to occur after the threshold is reached can vary from years to centuries [1]. Thus, an increase in the probability of the tipped state within the model signals that the real world system may be approaching a tipping point and that the probability of the system tipping in the future is increasing. Given that there is only one Earth, this seems to be ”too risky to bet against” [22].
@@ 7 Summary and Outlook
In this paper, we critically discussed potential pitfalls in visualizations of nonlinear dynamics used in both science communication and research. We discussed the need to distinguish between the author’s intended purpose of a visualization and the individual mental model that might be induced in the recipient in several examples, in particular for the ”effective potential” (see Fig. 3) which does not distinguish B- and N-tipping in the visualization. A key contribution of our analysis is the identification of crucial distinctions: the difference between deterministic and stochastic models, and between interpretations of a single system versus an ensemble of systems. Using examples representative of visualizations in popular and academic literature, we have shown that these distinctions are not always made. We therefore propose the visualizations shown in Fig. 5 for nonlinear dynamics and discuss examples from molecular biology (cell dynamics, Fig. 9, ensemble interpretation) and climate science (Amazon tipping point, Fig. 11, individual system) as explicit applications.
The visualizations we propose are not intended as a final solution, but rather as a theoretically grounded starting point for improved science communication. We acknowledge that the ultimate effectiveness of any visualization can only be validated through empirical research, which, however, requires well-defined and conceptually clear visualizations to test. By explicitly encoding the distinctions between the different cases discussed in 1, we believe our work provides this necessary starting point. The logical next step, therefore, is to employ empirical methods from physics education research and science communication to investigate the mental models that different audiences – from students to research professionals – construct from these proposed visualizations. This future research will be essential to test whether our refined representations indeed lead to a more accurate understanding of tipping point dynamics and help to avoid the potential pitfalls identified in this paper.